an account with $545 is invested at 5% simple interest for 6 years. what is the final balance in the account?

Found 2 solutions by rfer, shafqatlaghari:
Answer by rfer(16322)

I=545*0.05*6
I=163.5
-------
545+163.50=708.50

Answer by shafqatlaghari(2)

it is continuous compounding for 6 years so that
545 . (1.05)^6= 730.35